A new drug cures 9 of 200 patients suffering from a type of cancer, for which the historical cure rate is 2%. Perform a test to check on the significance of this result, at both 5% and 1% levels of significance. Based on your conclusion, comments on the efficacy of the treatment using this new drug. 


1
Expert's answer
2021-04-15T07:39:43-0400

"H_0 : p = 0.02 \\

H_1 : p ≠ 0.02 \\

n = 200 \\

\bar{p} = \frac{9}{200} = 0.045 \\

α_1 = 0.05"

Test statistic:

"Z = \frac{ \bar{p} -p}{\sqrt{ \frac{p(1-p)}{n}}} \\

Z = \frac{ 0.045 -0.02}{\sqrt{ \frac{(0.02)(0.08)}{200}}} = 2.53"

Critical value:

"Φ(Z_{cr}) = \frac{1 -α_1}{2} = 0.475 \\

Z_{cr}=1.96"

(−∞,−1.96) "\cup" (1.96,∞) is the rejection region.

The test statistic Z falls into the rejection region. So we reject H0 and accept H1.

The efficacy of the treatment using this new drug is not equal to 2%.

"α_2 = 0.01 \\

Φ(Z_{cr}) = \frac{1 -α_2}{2} = 0.495 \\

Z_{cr}=2.58"

(−∞,−2.58) "\cup" (2.58,∞) is the rejection region.

The test statistic Z does not fall into the rejection region. So we accept H0.

The efficacy of the treatment using this new drug is equal to 2%.